### v1

# Saccade crossing avoidance as a visual search strategy

Abstract
Although visual search appears largely random, several oculomotor biases exist such that the likelihoods of saccade directions and lengths depend on the previous scan path. Compared to the most recent fixations, the impact of the longer path history is more difficult to quantify. Using the step-selection framework commonly used in movement ecology, and analyzing data from 45-second viewings of Where’s Waldo?, we report a new memory-dependent effect that also varies significantly between individuals, which we term self-crossing avoidance. This is a tendency for saccades to avoid crossing those earlier in the scan path, and is most evident when both have small amplitudes. We show this by comparing real data to synthetic data generated from a memoryless approximation of the spatial statistics (i.e. a Markovian nonparametric model with a matching distribution of saccade lengths over time). Maximum likelihood fitting indicates that this effect is strongest when including the last $\approx 7$ seconds of a scan path. The effect size is comparable to well-known forms of history dependence such as inhibition of return. A parametric probabilistic model including a self-crossing penalty term was able to reproduce joint statistics of saccade lengths and self-crossings. We also quantified individual strategic differences, and their consistency over the six images viewed per participant, using mixed-effect regressions. Participants with a higher tendency to avoid crossings displayed smaller saccade lengths and shorter fixation durations on average, but did not display more horizontal, vertical, forward or reverse saccades. Together, these results indicate that the avoidance of crossings is a local orienting strategy that facilitates and complements inhibition of return, and hence exploration of visual scenes.
Statistically speaking, visual scan paths when viewing scenes look remarkably like random walks, particularly when salient areas are evenly distributed [1, 2]. Comparisons have frequently been made to animal foraging [3] — although in the case of vision, movement occurs primarily in discrete steps (saccades), between which are relatively still periods (fixations). Yet, as with animal foraging, there are task-dependent biases and patterns that can be uncovered if enough eye-tracking data is recorded [4].
Some biases, such as a tendency for the eyes to relax towards the centre of vision, can be specified as a simple function of the most recent fixation (i.e. fixations from this point toward the centre become more likely than in the opposite direction). In general, a Markovian model such as $P(x_{t+1},y_{t+1}|x_{t},y_{t})$ [5], when combined with saliency modelling, can come close to deep learning models in predicting the next fixation [6]. Such a distribution is also low-dimensional enough to generate with nonparametric methods [7].
Other biases depend on the most recent saccade (i.e. the two most recent fixations), such as the increased likelihood of forward and return saccades, which continue with the same length and in the same or reverse direction, respectively [8]. In other directions, regular changes in saccade length are more common than what is expected if subsequent lengths are assumed to be independent, leading to a negative autocorrelation [9].
Finally, some biases have longer history-dependence. The probability of saccades returning to ‘$k$-back’ fixations decreases only slowly with $k$ in some tasks [8]. This is has been called “facilitation of return” [10], in competition with “inhibition of return” which supposedly facilitates exploration [11]. It has been proposed that the former bias applies for salient points which require further consideration [8], while the latter is more prominent for simple distractors and low working memory demands [12].
The aforementioned influences on saccade dynamics have been quantified and modelled in various ways. Often they are based on comparing distributions of saccade and fixation measures to models assuming the absence and/or presence of some effect. The distributions compared (such as saccade length or change in saccade angles) depends on the phenomenon being examined. While there have been recent attempts to unify modeling approaches under the banner of likelihood-based inference [13], this has mainly been applied to comparing overall models against a benchmark rather than comparing the importance of predictors within models.
Another difficulty with identifying history dependence is the short length of most tasks and datasets. Deep learning models have been used to show that up to four fixations can increase predictive power in free viewing [14]. However, given what is known about refixations, it is expected that humans should be able to keep track of a longer scan path during search.
Figure 1: (a) List of predictors used (left), and an example of a scan path illustrating key quantities and case-control sampling (right). Predictors are sorted by the number of most recent fixations $L_{hist}$ required to calculate them. In the illustration, predictors are calculated for fixation $Q_{t+1}$ by generating four random control saccades (light dotted lines) in addition to the saccade from the data (dark dotted line). The center of the screen is indicated by a cross. (b) The panels on the left illustrate parametric predictor values for test points $Q^{\prime}_{t+1}$ given a scan path up to $Q_{t}$ indicated by a cross. The panels on the right show overall spatial log likelihoods for the nonparameteric model (top) the combined parameteric model (middle) and the difference (bottom).
Here, we use step selection analysis, a case-control sampling based method common in studies of animal movement [15], to quantify biases and long-term history dependence in scan paths during visual search. There are several advantages of this approach: Firstly, by using a typical saccade model to sample both taken and available steps, it can account for the fact that no two data points will have the same path history (in analogy with the heterogeneity of environmental features encountered by animals as they move). Secondly, being compatible with general regression models, it can evaluate the relative importance of and interactions between known biases in a straightforward way. Thirdly, by including random effects, it can also capture repeatable variation in biases between individuals [16], which would otherwise require a large amount of data per individual to quantify. Previous studies of strategic variation have focused on simple measures such as saccade frequency [17]. Given that visual search performance is correlated with several personality traits and cognitive abilities [18], we expect the underlying strategy to vary consistently along several other measures. While consistent differences have been identified in saccade lengths and fixation durations across tasks and time [19, 20, 21], stable patterns involving several measures have not yet been identified in visual search.
In addition to the most commonly reported biases and memory-dependent effects, we propose a new bias that we term self-crossing avoidance. This is inspired by the concept of space-filling curves, which exhibit this feature as a means to efficiently cover an area. The concept of self-avoiding random walks has also been explored by Engbert et al. to explain the statistics of small-scale eye movements during fixations [22], although in this case self-avoidance emerges from a combination of inhibition and a uniform microsaccade length. Instead, we include the number of recent saccades that are crossed by a candidate saccade as a potential influence on its likelihood.
We use the book series Where’s Waldo, which has been used for several other studies of visual search [11, 23, 24, 25]. With a high density of salient and heterogeneous distractors, this is a relatively difficult search task compared to natural scenes [25] and categorical target-distractor paradigms [26]. It is known that the density of visual features influence search patterns [26]. However, the relative homogeneity of the Waldo paradigm presents a task where the search space is not biased by more and less salient regions. Therefore, this is highly suitable for “pure” image-independent scan-path modelling, as well as for uncovering variation in search strategies.
To compare the search task to a bottom-up free viewing, we also use stimuli consisting of random coloured pixels, in which the participant is instructed to look for shapes and patterns hidden in the image. This presents another homogeneous stimulus favourable to identifying differences between individual scan patterns [27].
Results
The predictors used are defined in Figure 1 and visualized for a sample scan path. The predictors that depend on more than one previous saccade each contain a time constant that truncates the history length. These were optimized first separately and then jointly to maximize the likelihood of the logistic regression model (see Supplementary Information). For the Waldo task, the binary visited point predictor, indicating whether or not the candidate point was within 1.5 degrees (approximately the foveal angle) of a recent fixation, was optimal at $\tau_{1}=6$ s. The sparsity predictor, quantifying the ratio of the mean to maximum distance of previously visited points, was strongest when $\tau_{2}$ was set to 32 seconds. Finally, the self-crossing predictor, indicating the number of times a candidate saccade crossed the recent scan path, was optimal at $\tau_{3}=7$ seconds. When optimized jointly, $\tau_{1}$, $\tau_{2}$ and $\tau_{3}$ moved to 4, 33 and 6 seconds respectively. The three features therefore appear to capture complementary aspects of the path dependence, as shown by the lack of strong dependence between the time constants (see Supplementary Figure S1).
The sparsity measure, along with the distance from the center, partially capture how fixations were distributed overall within the screen area and among previous fixations. As shown in Figure 2(a-b), these distributions differ strongly between the the two tasks. Panel (b) shows that for the Waldo task, the likelihood of shifts in these quantities were more dependent on their current values; while (c) shows that, overall, changes in these predictors had opposite effects on fixation likelihood.
As expected, the history-independent predictors were the strongest overall, as shown in Figure 2(c) and Supplementary Figure S4. Parallel bias, which equally captures forward and return saccades, was the strongest linear predictor in the Waldo dataset. This and the additional forward bias were the only predictors that were in the same direction for the random pixel data. The predictors related to previously visited points indicated an inhibition of return on average in the Waldo task, as shown by the positive effect of change in sparsity and negative effect of a candidate fixation being close to one in its recent history, and conversely a facilitation of return for the random pixel task.
Figure 2: Task comparison. The panels in (a) compare the spatial distributions of fixations in terms of sparsity and distance from the centre of the image. Overlaid are moving averages over time, with a window length of 5 seconds, ending in the red crosses. The arrows in (b) shows the direction and slope of maximum gradient in log-odds from fitted parametric models evaluated at the small-saccade limit $Q^{\prime}_{t+1}\approx Q_{t}$ (see Supplementary Information). Panel (c) compares normalized z-scores for the main predictors in the model. The color of each cell gives the normalized z-score when including only a linear term for this predictor and all interactions for all other terms. Terms that do not decrease the AIC are given a normalized z-score of zero. Log-length, horizontal bias and cardinality are not included due to strong dependence on the saccade model.
Figure 3: (a) Mean number of crossings (including intersections with both past and future saccades within $\tau_{3}$) divided by saccade length, plotted against saccade length, for the models and test data. The grey area shows the density of saccade lengths in the test data, while the grey line shows the density of saccades with crossings. (b) Cumulative number of crossings over time for data generated from each model along with the test data. Shown are means and standard deviations over all 70 trials in each set.
For the Waldo task, synthetic data was generated to further investigate the self-crossing statistics. A nonparametric and memoryless null model was fit, based on a kernel-smoothed empirical distribution $P(\Delta x_{t},\Delta y_{t}|x_{t},y_{t})$. To control for the well-known coarse-to-fine transition [28], in which mean saccade length decreases over time, a cubic spline was fit to the saccade log-length as a function of time. Generated saccades were then rescaled according to this curve. The resulting generative model matched the overall distribution of saccade lengths for the entire trial length (see Figure 3(a) and Supplementary Figure S3). For comparison, the same amount of data was also generated using the parametric model fit by the step selection process.
For disconnected and independent line segments, crossing probability is expected to increase proportionally to length [29]. Here, since fixations are spatially clustered, there is a decreased probability of crossing per unit length for longer saccades (see Figure 3a). Compared to the null model, the experimental data showed relatively fewer crossings for all saccades, apart from the longest $\approx$ 15% (above 350 pixels) and those within the foveal angle. The trend was closely matched by the memory-dependent parametric model. This was also true for “crossing” and “crossed” saccades separately (see Supplementary Figure S5), despite existing saccade lengths not being a model input.
Overall, the mean number of crossings per saccade was $0.59$ compared to $0.33$. As shown in Figure 3b, the full parameteric model was closer to the crossing distribution of the data than of the null model, with a mean of $0.43$ crossings per saccade.
Figure 4: Waldo task individual differences. (a) Intraclass correlation coefficients for predictors where there was significant variation among both trials and participants. (b) Corresponding standard deviations of the participant-level random effect. (c) Pearson correlation coefficients between each significant random participant effect $(y)$ and the mean of the trial-level median fixation duration $t_{fix}$. Hatched bars indicate lack of significance at the $p<0.05$ level. (d) Loadings of the first principal component given the fitted means for each participant and each significant predictor.
Random effects were then added sequentially to each feature of a linear step selection model to quantify the individual variation in search strategies. For the Waldo task, all features had a significant participant effect (according to a lower AIC) apart from center distance change, fixation sparsity and fixation sparsity change. The intraclass-correlation coefficient, quantifying how much of the variation across trials was due to inter-participant variation, was highest for length change and self-crossing variance (ICC $0.91$ and $0.80$ respectively, see Figure 4(a)).
Several predictors were also positively correlated with those fitted for the random pixel task (without trial effects in the latter models due to the low number of trials per participant) at the $p<0.05$ level, with only cardinality, length change and visited predictors not being significant. The remaining correlations ranged from $r=0.42$ for self-crossing avoidance to $r=0.58$ for parallel bias.
The significant participant-level random effects for log length and self-crossings were positively correlated with average fixation durations in the same task (see Figure 4c, Supplementary Figure S8). When averaged over all participants, fixation time was greater for the random pixel task (mean 0.40 vs 0.27 for Waldo, t-test $t=10.3,p=10^{-11}$). Despite this large overall difference, the individual-level average fixation times were positively correlated between the two tasks (Pearson $r=0.43,p=0.016$).
When dimensionality of the unscaled random effects was reduced using principal component analysis, the first component captured 65% of the variance for the Waldo task. Participants that scored high on this component took longer steps, with more self-crossings (see Figure 4d). As was the case for these predictors on their own, the first principal component was also correlated with mean fixation duration ($r=0.54,p=0.002,n=31$). The second principal component, with 19% variance explained, was highest for participants with more forward bias, more horizontal bias, and less alternation in saccade length (see Supplementary Figure S9). The PCA loadings for the random pixel task were markedly different, and were not correlated with fixation time. However, the random effects for change in centre distance and parallel bias were negatively correlated with mean fixation time in this case (see Supplementary Figure S6).
Discussion
Difficult visual search tasks can take a considerable amount of time, providing ample opportunities to investigate participants’ search strategies as well as cognitive and oculomotor biases. The latter are more often studied using short-length free-viewing datasets designed for studying saliency, since these are much more widely available [6]. This means that long-term effects are generally under-appreciated when studying scan paths.
Using the step selection framework allowed various influences on the scan path to be considered simultaneously and to be compared on an equal footing. This technique is therefore suitable for tasks such as free viewing or visual search of scenes, and lessens the need for exhaustive experimental manipulations which, while often cleverly designed [30], ultimately lessen ecological validity. Although we did not use image features, assuming approximately uniform salience, step selection also allows the integration of aspects of saliency for comparing top-down and bottom-up influences on the scan path.
As is well known, participants looking at random, information-poor images will tend to focus more on the center of the image [27]. We have shown that in addition to the image center, several other influences have not only different strengths when comparing search and free viewing, but the effects are in opposite directions.
Since the resolution of spatial information in memory decreases over time [31], the longest term predictors should be the ones that require the least spatial resolution. Consistent with this, our fixation sparsity measure is insensitive to small changes in the scan path history, while the visited point predictor requires a resolution at the level of the foveal area.
The self-crossing tendency showed strong inter-individual variation and, on average, indicated avoidance during search. Those who avoided crossings more strongly also tended to make shorter saccades. This could mean that some participants pursued a systematic strategy of organized scan paths, as shown in [32] for stimuli on a grid. However, this would also imply strong variation in and correlation with cardinality, which was not seen in our data. We found that crossing shorter saccades tended to be avoided compared to the null model. This is consistent with the shorter optimal history length for this predictor.
Another clue for the functional significance of this variation can be found in the duration of fixations, which was absent from the step selection modelling. Fixation durations, widely used as a proxy for focused visual processing, were correlated at the individual level with saccade length and self-crossings in the Waldo task, which comprise the primary axis of variation in this task. Unlike the commonly found negative relationship between saccade length and fixation duration, as seen when comparing ambient and focal viewing over time [33], or between novices and experts [34], this relationship was instead found to be positive at the inter-individual level for the Waldo task. Such a relationship at the inter-individual level has been reported for free viewing [35], but previous studies for search, such as Henderson and Luke [21], may have had too small a sample size to reach statistical significance. Resolution of this effect may have also been enhanced in our case by using log-transforms for saccade length, which typically has a long-tailed distribution.
We also found that participants that were more likely to avoid self-crossings used shorter fixations on average. This may indicate that self-crossing avoidance is used primarily during the faster process of spatial orientation [36], and some participants use more of this “ambient” processing. In aviation, for example, experts use more complex scan paths and shorter fixations than amateurs [37]. Together, this may indicate that avoiding crossing of recent saccades is a relatively simple orienting strategy that facilitates exploration and complements explicit inhibition of return to recently visited locations.
This work also has implications for deep learning scan path models. The state-of-the-art Deepgaze III model [14] uses three images per fixation to encode their positions, so that they can be input into a convolutional neural network (CNN) and combined with a saliency map. This means that the size of the scanpath network will be proportional to the desired history length. Therefore, for deep learning to be useful in modeling human attention during complex search, it may be necessary to find more compact representations of history. Heatmaps of our three long time-scale heuristics, which we included for illustration, may also be useful as CNN inputs in this respect.
We have demonstrated a new and powerful method for quantifying oculomotor biases, leveraging techniques used in movement ecology. We expect that incorporating image features may provide both competitive and insightful results on common saliency benchmarks. Future work that correlates strategic variation with cognitive and personality differences may also shed light on the ways neural circuitry is recruited during visual search.
Materials and Methods
Data
31 participants, using chin rests approximately 1 meter from a 1920x1080 pixel monitor, were each shown nine Where’s Waldo images that filled the screen for 45 seconds each, followed by three images comprised of randomized RGB values per pixel, for 60 seconds each. Eye gaze was tracked using the EyeLink Duo at 1000Hz frame rate. Saccades and fixations were automatically extracted using the software. One in every four trials were reserved as a test set, distributed evenly among individuals and among images, with the remaining used for training the generative models. All data was used for the final fitting of mixed effect models.
Step selection
Importance sampling was used for generating control points [38, 39]. For each real data point, nine control data points were generated from a radially symmetric lognormal distribution fitted to the training set, with sampling repeated for any points generated outside the image boundary. For each data point, all predictors were calculated, along with the square of the log-length. These were then fed into a logistic regression to separate data points from control points. All interaction terms were initially included, and stepwise selection was used to remove insignificant interaction terms (see Supplementary Information). An interaction term between log length and time was also included to capture the coarse-to-fine transition [28]. These regressions were done using the Python statsmodels package.
Time constant optimization
For each variable-history predictor, the cutoff $\tau$ was iterated in units of one second up to the trial length of 45 seconds (see Supplementary Information). The value of $\tau$ maximizing the likelihood of the step selection, when including this predictor, the short-term predictors and all interactions, was then obtained and used for the rest of the analysis. For fixation sparsity, both the sparsity and the change in sparsity were included with the same value of $\tau_{2}$.
Nonparametric generative model
Gaussian kernel density estimation was used to smooth the conditional empirical distribution of $P(\Delta x_{i},\Delta y_{i}|x_{i},y_{i})$. One variance was used for $x,y$ and another was used for $\Delta_{x},\Delta,y$, and these were optimized jointly (see Supplementary Information). This was sampled from directly to generate synthetic data. To account for the change in average saccade length over time, a cubic spline was fit to the empirical log-length against time. Saccade lengths were then adjusted based on the time in the simulation according to the spline. Any points outside the boundary of the image were discarded and resampled. Time intervals between fixations were sampled from its unconditional distribution, which determined the total number of fixations within the trial length.
Parametric generative model
At each time step, 40 candidate points were generated as for the step selection process, resampling any outside the image boundary. The fitted step selection model was then used to estimate the likelihood of each point, adjusting the intercept term for the difference in number of samples. One sample was then chosen with a probability proportional to this likelihood. Time intervals were sampled from the unconditional distribution as for the nonparametric model.
Mixed-effect modelling
Individual variation was tested and quantified using a binomial mixed effect regression. All predictors were first centered and normalized. In the Waldo task, for each of the fixed effects, a term modeling the slope varying over trials (279, nine per participant, variance $\sigma_{t}^{2}$) was tested first against the base model with only fixed effects using the AIC. If the mixed model had a lower AIC, a second random slope effect was included for participant (31, variance $\sigma_{p}^{2}$), and also tested using the AIC. For predictors with a significant participant effect, the ratio $\sigma_{p}^{2}/(\sigma_{p}^{2}+\sigma_{t}^{2})$ was used as the intraclass correlation coefficient (ICC). For the random pixel task, only one test was performed for the participant effect, due to the small number of trials, and thus the ICC was not obtained. The R package lme4 was used via the Python package pymer4[40].
Acknowledgements
This work is funded by the Research Council of Norway’s FRIPRO grant number 335940.